### v1

@@
1 Introduction
Foundation Models (FMs), particularly Large Language Models (LLMs), are transforming software engineering by enabling FMware, i.e., software systems that integrate FMs as core components. FMware expands software capabilities, allowing domain experts to develop AI-driven applications without extensive programming expertise. As adoption accelerates, the market for FMware is projected to grow at a compound annual rate of 35.9% through 2030 [llmmarket]. Despite this rapid growth, developing FMware remains a significant challenge, requiring engineering practices that differ fundamentally from traditional software development. Building FMware is difficult because it introduces stochastic behavior, evolving dependencies, and high computational costs while lacking established engineering methodologies. Developers must handle unpredictable outputs, data sensitivity, prompt brittleness, and limited interpretability, making even reliable prototyping difficult. FMware often integrates multiple AI components, creating interdependencies that amplify failure points and complicate engineering. These complexities set FMware apart from conventional software, demanding new software engineering approaches. Deploying FMware at scale presents an even greater challenge. Production-ready FMware must ensure performance stability, robustness, and compliance while managing scalability constraints, security risks, and deployment costs. Industry experiences illustrate these difficulties: LinkedIn required four additional months to refine the final 20% of an FMware deployment, facing diminishing returns on each incremental improvement [linkedinMusingsBuilding]. Microsoft and GitHub reported that as FMware complexity increases, validation and testing become prohibitively expensive [parnin2023building] due to the absence of standardized evaluation frameworks. The cost of deployment further complicates production readiness. For instance, OpenAI’s ChatGPT infrastructure alone was estimated to cost \$700,000 per day in 2023 [uniteFinancialChallenges]. Existing research has explored FMs’ role in software engineering, including their applications in code generation, testing, documentation, and reasoning [jin2024llms, fan2023large, zhang2023survey, hou2024large, jiang2024survey]. These studies primarily rely on academic benchmarks and structured datasets, often overlooking real-world deployment challenges. Meanwhile, research on software production readiness has largely focused on release criteria, system reliability, and ML-specific deployment challenges [asthana2009quantifying, cusick2013architecture, breck2017ml, parnin2023building, oreillyWhatLearned]. Studies on enterprise FM integration [nahar2024beyond, chen2025empirical] highlight developer concerns but remain constrained to specific ecosystems. A recent survey accompanying a KDD 2024 tutorial [ding2024reasoning] examined how LLMs enhance software development workflows. While that work focused on improving development processes using LLMs, we examine a different challenge: the engineering complexities of FMware itself. Unlike prior surveys, which emphasize how LLMs assist software engineering, we analyze the core software engineering difficulties of building, deploying, and maintaining FMware at scale. Furthermore, this survey improves on previous work [rajbahadur2025cooldemosproductionreadyfmware, hassan2024rethinking] by providing a cohesive overview of all challenges related to FMware development in a way that is relevant to researchers and practitioners alike. Our work systematically identifies these challenges and provides a roadmap for addressing them. To bridge these gaps, we propose a KDD 2025 lecture-style tutorial that explores insights, concerns, and practical strategies for developing and deploying trustworthy, production-ready FMware. We first present an overview of the FMware lifecycle in Section 2. We then discuss key the challenges and present a critical overview of the state of practice in FMware engineering in Section 3. Finally, we outline actionable insights and emerging solutions, offering a roadmap for advancing FMware engineering in Section 4. Section 5 concludes our paper.
2 FMware Lifecycle
[…]
[width=1.0]figs/fmware_lifecycle.pdf
[…]
@@ 2 FMware Lifecycle
[…]
Selecting an appropriate FM is the first step in the FMware lifecycle. Unlike traditional software components with predefined logic, FMs generate responses based on statistical patterns from large-scale training data. Developers must consider task requirements, scalability, licensing constraints, and adaptability when choosing an FM. Once selected, alignment is necessary to fine-tune an FM for application-specific needs. Data alignment involves curating domain-relevant datasets, ensuring the FM performs reliably on specialized tasks [liu2024best, albalak2024survey]. Model alignment further refines behavior using techniques such as instruction tuning and reinforcement learning with human feedback (RLHF) to guide model outputs toward desired behaviors [shen2023large, xia2024understanding]. {comment} State of Practice. To reduce manual effort in alignment, researchers explore active learning [yu2022actune, toloka] and weak supervision [biegel2021active, bach2019snorkel]. Approaches like Snorkel AI [ratner2017snorkel] and Self-Instruct [wang2023selfinstruct] generate labeled datasets automatically, though manual oversight is still required [maekawa2022low]. Recent studies focus on data quality over quantity, showing that fine-tuning on high-quality instructional data reduces alignment costs [gunasekar2023textbooks, zhou2023lima]. To improve debugging, researchers employ influence functions and feature importance analysis to identify impactful training data points [grosse2023studying, schoch2023data]. While traditional data testing tools [braiek2020testingml, krishnan2016activeclean] aid dataset cleaning, they struggle with the high dimensionality and opacity of FMs, making fine-grained debugging an ongoing challenge.
[…]
Since FMware interacts with users through natural language prompts rather than explicit function calls, prompt engineering defines system behavior. Developers iteratively refine prompts to improve accuracy, consistency, and adaptability across different contexts [parthasarathy2024ultimate]. Production-ready systems require structured prompt design, incorporating formatting, context augmentation, and constraint mechanisms. To enhance reliability, developers employ well-established prompt patterns, such as chain-of-thought prompting [chainOfThought2022], self-consistency prompting, and structured templating. These patterns improve coherence, reduce variability, and enable more predictable FM responses. Prompt portability is another critical consideration, ensuring that prompts remain effective when transitioning between different FMs [hassan2024rethinking]. However, manually crafting and optimizing prompts remains a labor-intensive and iterative process, leading researchers and practitioners to explore automated and systematic solutions. Grounding improves response reliability by linking FM outputs to external information sources. Techniques such as retrieval-augmented generation (RAG) ensure factual accuracy and reduce hallucination risks [lewis2020rag]. Grounding also aids compliance by ensuring alignment with authoritative sources, mitigating risks of generating misleading content [nahar2024beyond]. {comment} State of Practice. To address the inefficiencies in prompt development, various tools and methods have emerged. Visual tools like Prompt IDE [xPromptIDE] and Aleph Alpha [deiseroth2023atman] provide inference transparency and token-level explainability to assist FMware developers in refining prompts. However, they lack higher-level debugging capabilities for analyzing logical inconsistencies. Researchers explore semi-automated approaches [hsieh2023automatic, xu2022gps, ye2023prompt] and reinforcement learning-based optimization [deng2022rlprompt] to improve prompt effectiveness. For example, Automatic Prompt Engineer [zhou2023large] and PromptBreeder [fernando2023promptbreeder] use evolutionary search techniques to generate optimal prompts, while DSPy [khattab2023dspy] integrates prompt optimization with a programming model. However, these techniques often optimize prompts in isolation, ignoring variations in model candidates, inference parameters, and hardware environments, which are critical in real-world FMware deployment.
[…]
While FMware can operate through direct prompt-response interactions, modern applications increasingly incorporate autonomous agents for decision-making, task execution, and interoperability with external systems. Agents extend FM capabilities by enabling multi-step reasoning, maintaining state, and integrating with traditional software components. An FMware agent functions as an adaptive entity that executes complex workflows, interacts with external tools and databases, and refines decision-making through feedback loops. To manage these capabilities, cognitive architecture planning defines how FMware components coordinate within a system and enable agent orchestration. Depending on the application, orchestration can take different forms: static workflows, where execution paths are predefined; tool-augmented execution, where FMs interact with databases or predefined functions dynamically; and agent-based orchestration, where autonomous agents adjust execution flows and collaborate with other agents in real time. FMware agents integrate with traditional software architectures, enhancing existing deterministic systems by bridging structured programming with adaptive reasoning. {comment} State of Practice. To facilitate multi-generational software integration, current practices encapsulate legacy systems as tools that FMware can access. Frameworks like Semantic Kernel [semantickernel] use plugins and native functions to abstract legacy system capabilities into modular components. AutoGen [wu2023autogen] and Copilot Workspace [githubnextcopilotworkspace] offer mechanisms for human oversight, allowing users to guide agent behavior. MetaGPT [hong2023metagpt] encodes standard operating procedures (SOPs) into prompts, while symbolic planning approaches [zhou2023agents] enhance controllability in agent-driven systems. These methods highlight the importance of balancing automation with human oversight to ensure reliable FMware orchestration.
[…]
Observability in software engineering refers to the ability to monitor system behavior, detect failures, and analyze performance trends. Traditional observability relies on logs, traces, and metrics to ensure expected functionality. However, FMware introduces new challenges due to its probabilistic nature, requiring deeper insights into dynamic behaviors beyond conventional monitoring. Semantic observability extends traditional practices by capturing user corrections, intent recognition, and model drift detection to refine FMware performance over time. Unlike classical observability, which tracks system health at a functional level, FMware observability must also analyze emergent behaviors arising from model interactions and decision-making processes. Ensuring trust in FMware requires robust guarding mechanisms to enforce compliance, safety, and reliability [rebedea2023nemo, dong2024framework]. Unlike static rule-based validation, FMware must assess correctness dynamically at inference time, incorporating safeguards for data validation, response filtering, and risk mitigation. {comment} State of Practice. Efforts to enhance FMware observability are advancing but remain largely adapted from traditional software monitoring. Tools like OpenLLMetry [traceloopOpensourceObservability] and LangSmith [langchainLangSmith] capture traces of FM calls, vector database interactions, and latency metrics. However, they lack insights into the cognitive processes underlying FMware decisions, especially in multi-agent environments where behaviors emerge dynamically. GitHub Copilot employs semantic signal telemetry by tracking user edits after code generation, but generalizable solutions for FMware-wide semantic monitoring are still lacking. More sophisticated frameworks, such as Microsoft PowerAutomate, offer workflow-level tracing, allowing developers to diagnose agent coordination issues. However, these approaches focus on functional observability rather than capturing how agents plan, reason, and interact. While LangSmith provides structured logging for LangChain-based applications, it remains limited to single-agent workflows, leaving multi-agent observability an open challenge.
[…]
FMware validation differs from traditional software testing due to its non-deterministic outputs. Unlike conventional applications, where the same input produces the same output, FMware responses vary based on model updates, prompt changes, and shifting input distributions [ma2024my]. Ensuring reliability requires continuous evaluation pipelines that benchmark consistency across evolving model versions while assessing distributional stability, response variability, and edge case handling [microsoftRetryStorm]. Optimization efforts extend beyond accuracy, balancing latency, compute efficiency, and cost constraints. Unlike deterministic APIs, FM inference introduces variable processing times, requiring adaptive strategies such as query batching, caching, and model routing. {comment} State of Practice. FMware requires scalable solutions that span multiple models and inference rounds. To mitigate performance issues, developers prioritize smaller models with lower latency. For example, Meta selected the 1.3B InCoder model over a 6.7B alternative for code authoring, prioritizing speed over minor accuracy trade-offs [murali2023codecompose]. Compression techniques such as quantization, distillation, and pruning [zhu2023survey] improve inference speed, but balancing efficiency without degrading performance remains an open challenge. Prompt compression [jiang2023llmlingua] reduces input size but risks losing key information, while semantic caching [gim2023prompt] accelerates responses by reusing previously generated outputs. Current FMware evaluation heavily relies on human effort or automated validation using high-performing models such as GPT-4 [liu2023gpteval]. However, FM-based evaluators only achieve moderate correlation (0.51) with human assessments, limiting their reliability [liu2023gpteval]. AI-as-judge methods are typically widely used, FM-based evaluators are known to favor superficial attributes over factual accuracy, resulting in inconsistent judgments [thakur2024judging, chen2024humans]. Traditional software testing frameworks fail to address FMware’s resource constraints and dynamic nature. Existing caching solutions, such as those used in LangChain, do not account for real-time updates and prompt variations, leading to ineffective cache invalidation. Additionally, FMware lacks dependency-based test execution [sharif2021deeporder], resulting in redundant tests that increase computational costs. Retry mechanisms remain underdeveloped, as FMware failures often require adaptive re-prompting instead of simple re-execution, necessitating more intelligent fault-tolerance strategies.
[…]
FMware requires efficient memory management to preserve context across interactions, optimize multi-agent workflows, and manage inference workloads [microsoftRetryStorm, luo2024arena]. Unlike traditional applications with explicitly stored state, FMware must dynamically handle context persistence, agent coordination, and compute optimization to ensure efficiency. {comment} State of Practice. Effective memory management remains a key challenge in FMware, particularly in synchronizing knowledge across multiple FMs and user contexts. Systems often struggle with consistency in multi-feedback loops [microsoftCheckYour], making feedback integration unreliable. During deployment, memory handling supports model swapping and prevents performance degradation from latency issues [llmpilot]. Managing retries efficiently avoids costly retry storms while ensuring agents share state and context for coherent decision-making [microsoftRetryStorm, luo2024arena]. Robust memory systems are also critical for regulatory compliance and optimizing dynamic workloads [hassan2024rethinking, wang2023survey]. Knowledge representation inefficiencies inflate computational costs and response times by storing redundant or irrelevant data. Poor retrieval optimization can lead to inconsistent reasoning and incorrect outputs [guo2024knowledgenavigator, wu2024easily]. Additionally, managing memory across FMware instances is error-prone due to synchronization issues, conflicting knowledge, and data corruption [xie2023adaptive, chen2022rich]. Production-ready FMware requires robust memory architectures to mitigate context loss and memory mismanagement, ensuring stable and reliable performance in deployment environments [packer2023memgpt].
[…]
Compliance in FMware spans the entire lifecycle, requiring continuous evaluation due to the probabilistic nature of model outputs and evolving regulations. Organizations must ensure that training data and model alignment comply with licensing, privacy laws, and ethical AI principles [hassan2024rethinking, nahar2024beyond]. Regulatory risks arise from data misuse, while model fine-tuning techniques such as RLHF must be audited to prevent biases. Inference-time compliance demands real-time safeguards, including content filtering, bias detection, and automated moderation to mitigate harmful outputs [dong2024framework, rebedea2023nemo]. Post-deployment, FMware requires robust logging, auditing, and transparency measures such as model cards and AI Bill of Materials (AI BOM) [hassan2024rethinking]. Governance frameworks must define accountability, risk management, and alignment with emerging AI regulations [nahar2024beyond], ensuring responsible FMware deployment while maintaining adaptability to regulatory shifts. {comment} State of Practice. Regulatory compliance in FMware remains largely manual and process-heavy. Efforts such as enhanced model and data cards [brajovic2023model] and AI risk management frameworks [nist2021guidance, novelli2023taking] provide guidance but lack scalability for complex FMware systems. License compliance has seen advances with Responsible AI Licenses (RAIL) [contractor2022behavioral] and OpenDataology [opendataology], but these solutions often focus on individual models or datasets rather than the integrated FMware ecosystem. Emerging standards like SPDX 3.0 Dataset Profile [SPDXDatasetProfile2023] and AI BOMs offer promising directions for automating compliance and tracking dependencies, though gaps remain in addressing compliance for synthetic data and evolving FMware components, particularly in scenarios involving agent-driven learning and adaptation. The FMware lifecycle defines how foundation models integrate into modern software systems, introducing new considerations for prompt engineering, agent orchestration, observability, and continuous adaptation. Unlike traditional software, FMware evolves dynamically through interaction feedback, semantic observability, and model-driven optimization, requiring engineering approaches that blend AI, software reliability, and real-time adaptability. As FMware adoption accelerates, structured software engineering principles will be critical to ensuring scalability, trustworthiness, and enterprise-level integration.
3 Development & Productionization Challenges in FMware

### v2

@@
Abstract.
Foundation Models (FMs) such as Large Language Models (LLMs) are reshaping the software industry by enabling FMware, systems that integrate these FMs as core components. In this KDD 2025 tutorial, we present a comprehensive exploration of FMware that combines a curated catalogue of challenges with real-world production concerns. We first discuss the state of research and practice in building FMware. We further examine the difficulties in selecting suitable models, aligning high-quality domain-specific data, engineering robust prompts, and orchestrating autonomous agents. We then address the complex journey from impressive demos to production-ready systems by outlining issues in system testing, optimization, deployment, and integration with legacy software. Drawing on our industrial experience and recent research in the area, we provide actionable insights and a technology roadmap for overcoming these challenges. Attendees will gain practical strategies to enable the creation of trustworthy FMware in the evolving technology landscape.
Foundation models, Large Language Models, AI4SE
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736572isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Natural language generationccs: Computing methodologies Artificial intelligenceccs: Software and its engineering Software development techniques
1. Introduction
Foundation Models (FMs), particularly Large Language Models (LLMs), are transforming software engineering by enabling FMware, i.e., software systems that integrate FMs as core components. FMware expands software capabilities, allowing domain experts to develop AI-driven applications without extensive programming expertise. As adoption accelerates, the market for FMware is projected to grow at a compound annual rate of 35.9% through 2030 (Grand View Research, 2023). Despite this rapid growth, developing FMware remains a significant challenge, requiring engineering practices that differ fundamentally from traditional software development.
Building FMware is difficult because it introduces stochastic behavior, evolving dependencies, and high computational costs while lacking established engineering methodologies. Developers must handle unpredictable outputs, data sensitivity, prompt brittleness, and limited interpretability, making even reliable prototyping difficult. FMware often integrates multiple AI components, creating interdependencies that amplify failure points and complicate engineering. These complexities set FMware apart from conventional software, demanding new software engineering approaches.
Deploying FMware at scale presents an even greater challenge. Production-ready FMware must ensure performance stability, robustness, and compliance while managing scalability constraints, security risks, and deployment costs. Industry experiences illustrate these difficulties: LinkedIn required four additional months to refine the final 20% of an FMware deployment, facing diminishing returns on each incremental improvement (lin, 2024). Microsoft and GitHub reported that as FMware complexity increases, validation and testing become prohibitively expensive (Parnin et al., 2023) due to the absence of standardized evaluation frameworks. The cost of deployment further complicates production readiness. For instance, OpenAI’s ChatGPT infrastructure alone was estimated to cost \$700,000 per day in 2023 (Abbas, 2024).
Existing research has explored FMs’ role in software engineering, including their applications in code generation, testing, documentation, and reasoning (Jin et al., 2024; Fan et al., 2023; Zhang et al., 2023; Hou et al., 2024; Jiang et al., 2024). These studies primarily rely on academic benchmarks and structured datasets, often overlooking real-world deployment challenges. Meanwhile, research on software production readiness has largely focused on release criteria, system reliability, and ML-specific deployment challenges (Asthana and Olivieri, 2009; Cusick, 2013; Breck et al., 2017; Parnin et al., 2023; Yan et al., [n. d.]). Studies on enterprise FM integration (Nahar et al., 2024; Chen et al., 2025) highlight developer concerns but remain constrained to specific ecosystems. A recent survey accompanying a KDD 2024 tutorial (Ding et al., 2024) examined how LLMs enhance software development workflows. While that work focused on improving development processes using LLMs, we examine a different challenge: the engineering complexities of FMware itself. Unlike prior surveys, which emphasize how LLMs assist software engineering, we analyze the core software engineering difficulties of building, deploying, and maintaining FMware at scale. Furthermore, this survey improves on previous work (Rajbahadur et al., 2025; Hassan et al., 2024a) by providing a cohesive overview of all challenges related to FMware development in a way that is relevant to researchers and practitioners alike. Our work systematically identifies these challenges and provides a roadmap for addressing them.
To bridge these gaps, we propose a KDD 2025 lecture-style tutorial that explores insights, concerns, and practical strategies for developing and deploying trustworthy, production-ready FMware. We first present an overview of the FMware lifecycle in Section 2. We then discuss key the challenges and present a critical overview of the state of practice in FMware engineering in Section 3. Finally, we outline actionable insights and emerging solutions, offering a roadmap for advancing FMware engineering in Section 4. Section 5 concludes our paper.
2. FMware Lifecycle
Figure 1. FMware Lifecycle.
@@ 2. FMware Lifecycle
2.1. FM Selection and Alignment
Selecting an appropriate FM is the first step in the FMware lifecycle. Unlike traditional software components with predefined logic, FMs generate responses based on statistical patterns from large-scale training data. Developers must consider task requirements, scalability, licensing constraints, and adaptability when choosing an FM.
Once selected, alignment is necessary to fine-tune an FM for application-specific needs. Data alignment involves curating domain-relevant datasets, ensuring the FM performs reliably on specialized tasks (Liu et al., 2024b; Albalak et al., 2024). Model alignment further refines behavior using techniques such as instruction tuning and reinforcement learning with human feedback (RLHF) to guide model outputs toward desired behaviors (Shen et al., 2023; Xia et al., 2024).
2.2. Prompt Engineering and Grounding
Since FMware interacts with users through natural language prompts rather than explicit function calls, prompt engineering defines system behavior. Developers iteratively refine prompts to improve accuracy, consistency, and adaptability across different contexts (Parthasarathy et al., 2024). Production-ready systems require structured prompt design, incorporating formatting, context augmentation, and constraint mechanisms.
To enhance reliability, developers employ well-established prompt patterns, such as chain-of-thought prompting (Wei et al., 2022), self-consistency prompting, and structured templating. These patterns improve coherence, reduce variability, and enable more predictable FM responses. Prompt portability is another critical consideration, ensuring that prompts remain effective when transitioning between different FMs (Hassan et al., 2024a). However, manually crafting and optimizing prompts remains a labor-intensive and iterative process, leading researchers and practitioners to explore automated and systematic solutions.
Grounding improves response reliability by linking FM outputs to external information sources. Techniques such as retrieval-augmented generation (RAG) ensure factual accuracy and reduce hallucination risks (Lewis et al., 2020). Grounding also aids compliance by ensuring alignment with authoritative sources, mitigating risks of generating misleading content (Nahar et al., 2024).
2.3. Agents, Orchestration, and Cognitive Architecture Planning
While FMware can operate through direct prompt-response interactions, modern applications increasingly incorporate autonomous agents for decision-making, task execution, and interoperability with external systems. Agents extend FM capabilities by enabling multi-step reasoning, maintaining state, and integrating with traditional software components. An FMware agent functions as an adaptive entity that executes complex workflows, interacts with external tools and databases, and refines decision-making through feedback loops.
To manage these capabilities, cognitive architecture planning defines how FMware components coordinate within a system and enable agent orchestration. Depending on the application, orchestration can take different forms: static workflows, where execution paths are predefined; tool-augmented execution, where FMs interact with databases or predefined functions dynamically; and agent-based orchestration, where autonomous agents adjust execution flows and collaborate with other agents in real time.
FMware agents integrate with traditional software architectures, enhancing existing deterministic systems by bridging structured programming with adaptive reasoning.
2.4. Observability and Guarding
Observability in software engineering refers to the ability to monitor system behavior, detect failures, and analyze performance trends. Traditional observability relies on logs, traces, and metrics to ensure expected functionality. However, FMware introduces new challenges due to its probabilistic nature, requiring deeper insights into dynamic behaviors beyond conventional monitoring.
Semantic observability extends traditional practices by capturing user corrections, intent recognition, and model drift detection to refine FMware performance over time. Unlike classical observability, which tracks system health at a functional level, FMware observability must also analyze emergent behaviors arising from model interactions and decision-making processes.
Ensuring trust in FMware requires robust guarding mechanisms to enforce compliance, safety, and reliability (Rebedea et al., 2023; Dong et al., 2024). Unlike static rule-based validation, FMware must assess correctness dynamically at inference time, incorporating safeguards for data validation, response filtering, and risk mitigation.
2.5. Performance Engineering, Testing, and Optimization
FMware validation differs from traditional software testing due to its non-deterministic outputs. Unlike conventional applications, where the same input produces the same output, FMware responses vary based on model updates, prompt changes, and shifting input distributions (Ma et al., 2024). Ensuring reliability requires continuous evaluation pipelines that benchmark consistency across evolving model versions while assessing distributional stability, response variability, and edge case handling (RobBagby, [n. d.]). Optimization efforts extend beyond accuracy, balancing latency, compute efficiency, and cost constraints. Unlike deterministic APIs, FM inference introduces variable processing times, requiring adaptive strategies such as query batching, caching, and model routing.
2.6. Deployment, Maintenance, and Tooling
Deploying FMware extends beyond traditional CI/CD practices, as model behavior evolves dynamically rather than following static updates (Łazuka et al., 2024). Maintaining FMware requires continuous inference monitoring, prompt tuning based on observed performance, and version rollback mechanisms to mitigate performance regressions. Unlike conventional software, FMware lacks standardized tooling for versioning, fine-tuning management, and scalable inference deployment, necessitating adaptive update strategies to ensure stability.
2.7. Cross-cutting concerns
2.7.1. Memory Management
FMware requires efficient memory management to preserve context across interactions, optimize multi-agent workflows, and manage inference workloads (RobBagby, [n. d.]; Luo et al., 2024). Unlike traditional applications with explicitly stored state, FMware must dynamically handle context persistence, agent coordination, and compute optimization to ensure efficiency.
2.7.2. Compliance & Regulations
Compliance in FMware spans the entire lifecycle, requiring continuous evaluation due to the probabilistic nature of model outputs and evolving regulations. Organizations must ensure that training data and model alignment comply with licensing, privacy laws, and ethical AI principles (Hassan et al., 2024a; Nahar et al., 2024). Regulatory risks arise from data misuse, while model fine-tuning techniques such as RLHF must be audited to prevent biases.
Inference-time compliance demands real-time safeguards, including content filtering, bias detection, and automated moderation to mitigate harmful outputs (Dong et al., 2024; Rebedea et al., 2023). Post-deployment, FMware requires robust logging, auditing, and transparency measures such as model cards and AI Bill of Materials (AI BOM) (Hassan et al., 2024a). Governance frameworks must define accountability, risk management, and alignment with emerging AI regulations (Nahar et al., 2024), ensuring responsible FMware deployment while maintaining adaptability to regulatory shifts.
The FMware lifecycle defines how foundation models integrate into modern software systems, introducing new considerations for prompt engineering, agent orchestration, observability, and continuous adaptation. Unlike traditional software, FMware evolves dynamically through interaction feedback, semantic observability, and model-driven optimization, requiring engineering approaches that blend AI, software reliability, and real-time adaptability. As FMware adoption accelerates, structured software engineering principles will be critical to ensuring scalability, trustworthiness, and enterprise-level integration.
3. Development & Productionization Challenges in FMware